# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 4. Koszul duality
$\tau_{\mathrm{n}}({\mathcal{O}})*_{{\mathcal{O}}}{\mathrm{triv}}\simeq\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}}.$
Construction 4.11.
@@
in ${\mathcal{C}}$ for any right ${\mathcal{O}}^{\vee}$-comodule ${\mathrm{Z}}$ in ${\mathrm{sSeq}}({\mathcal{C}}).$
Recall that by [23, Example 6.1.6.22.] for every object ${\mathrm{X}}$ of ${\mathcal{C}}$ equipped with an action of a finite group ${\mathrm{G}}$ there is a norm map ${\mathrm{X}}_{\mathrm{G}}\to{\mathrm{X}}^{\mathrm{G}}$ from the coinvariants to the invariants, which we will use in the following.
@@
Let ${\mathcal{C}}$ be a stable symmetric monoidal $\infty$-category compatible with small colimits that admits small limits, ${\mathcal{O}}$ a non-unital $\infty$-operad in ${\mathcal{C}}$ such that ${\mathcal{O}}_{1}={\mathbb{1}}$ and ${\mathrm{X}}$ an ${\mathcal{O}}$-algebra. If all norm maps associated to symmetric groups are equivalences, for every ${\mathrm{n}}\in{\mathbb{N}}$ the following morphism is an equivalence:
$\rho:\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}}\to(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{{\mathcal{O}}}{\mathrm{X}}).$
@@
We proceed by induction on ${\mathrm{n}}\geq 1.$ For ${\mathrm{n}}=1$ the map $\rho$ identifies with the canonical equivalence
@@
The fiber ${\mathrm{T}}$ of the map $\tau_{{\mathrm{n}}+1}({\mathcal{O}})\to\tau_{\mathrm{n}}({\mathcal{O}})$ of right ${\mathcal{O}}$-modules is the symmetric sequence concentrated in degree ${\mathrm{n}}+1$ with value ${\mathcal{O}}_{{\mathrm{n}}+1}$ equipped with the trivial right ${\mathcal{O}}$-module structure according to Lemma 2.46. Applying the exact functor $(-)\circ_{\mathcal{O}}{\mathrm{X}}:{\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}}))\to{\mathrm{sSeq}}({\mathcal{C}})$ we obtain the fiber sequence
${\mathrm{T}}\circ{\mathrm{Y}}={\mathrm{T}}\circ({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})\to\tau_{{\mathrm{n}}+1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}\to\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}.$
@@
$(\tau_{{\mathrm{n}}+1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}){*}^{{\mathcal{O}}^{\vee}}{\mathrm{Y}}\to(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}){*}^{{\mathcal{O}}^{\vee}}{\mathrm{Y}}$
@@
${{\mathrm{T}}\circ{\mathrm{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tau_{{\mathrm{n}}+1}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{T}}{\ast}{\mathrm{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\tau_{{\mathrm{n}}+1}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}}){\ast}^{{\mathcal{O}}^{\vee}}{\mathrm{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}}){\ast}^{{\mathcal{O}}^{\vee}}{\mathrm{Y}},}$
where both horizontal sequences are fiber sequences. The left vertical map ${\mathrm{T}}\circ{\mathrm{Y}}\to{\mathrm{T}}{\ast}{\mathrm{Y}}$ identifies with the norm map $({\mathcal{O}}_{{\mathrm{n}}+1}\otimes{\mathrm{Y}}^{\otimes{\mathrm{n}}+1})_{\Sigma_{{\mathrm{n}}+1}}\to({\mathcal{O}}_{{\mathrm{n}}+1}\otimes{\mathrm{Y}}^{\otimes{\mathrm{n}}+1})^{\Sigma_{{\mathrm{n}}+1}}.$
@@
${{\mathrm{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{triv}}*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\lim_{{\mathrm{n}}\geq 1}(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\lim_{{\mathrm{n}}\geq 1}((\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})).}$
@@
${{\mathrm{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{triv}}*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\lim_{{\mathrm{n}}\geq 1}(\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\lim_{{\mathrm{n}}\geq 1}((\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})).}$
@@
At the time of writing this article we learned about an unpublished article [15] that generalizes Proposition 4.13 to the situation that norm maps are not necessarily invertible.
@@ 5. A derived Milnor-Moore theorem
Computing the Koszul dual one finds that for every ${\mathrm{n}}\in\Sigma$ there is a canonical equivalence of spectra
${\mathrm{Lie}}_{\mathrm{n}}\simeq\bigoplus_{{\mathrm{n}}-1!}{\mathbb{1}}_{\mathrm{Sp}},$
while the $\Sigma_{\mathrm{n}}$-action on ${\mathrm{Lie}}_{\mathrm{n}}$ is highly non-trivial. In particular, ${\mathrm{Lie}}$ is an $\infty$-operad in connective spectra $\mathrm{Sp}_{\geq 0}.$
@@
${\mathrm{Lie}}_{\mathcal{C}}=\alpha^{\mathcal{C}}_{!}({\mathrm{Lie}})\simeq({\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}})^{\vee}(-1).$
@@
    $\theta:{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
    ${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})$
@@
1. We like to see that for any ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})$ the canonical map $\theta({\mathrm{X}}\times{\mathrm{Y}})\to\theta({\mathrm{X}})\times\theta({\mathrm{Y}})$ is an equivalence. By Lemma 2.30 the forgetful functors ${\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathcal{C}}$ and ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathcal{C}}$ are comonadic. Let $\mathrm{Sym}_{>0}$ be the right adjoint of the forgetful functor ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathcal{C}}$. By comonadicity any ${\mathrm{X}}\in{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})$ is the totalization of a cosimplicial object ${\mathrm{X}}^{\prime}$ in ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})$ taking values in the essential image of $\mathrm{Sym}_{>0}$, whose image in ${\mathcal{C}}$ splits. Hence $\theta$ preserves the limit of ${\mathrm{X}}^{\prime}$ and the canonical map $\theta({\mathrm{X}}\times{\mathrm{Y}})\to\theta({\mathrm{X}})\times\theta({\mathrm{Y}})$ identifies with the map
@@
preserves finite products. By [23, Proposition 3.2.4.7.] the object-wise symmetric monoidal structure on ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ is cartesian. For any ${\mathrm{X}},{\mathrm{Y}}\in{\mathcal{C}}$ the canonical map $\mathrm{Sym}({\mathrm{X}}\oplus{\mathrm{Y}})\to\mathrm{Sym}({\mathrm{X}})\otimes\mathrm{Sym}({\mathrm{Y}})$ identifies with the canonical equivalence in ${\mathcal{C}}.$
2.: In view of (1) it is enough to see that the composition
$\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
preserves finite products.
In the following we consider ${\mathbb{N}},{\mathbb{Z}}$ as categories by viewing them as posets. Let
$\gamma:\mathrm{Fun}({\mathbb{Z}},{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\to\prod_{{\mathbb{Z}}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
be the functor that sends a filtered object ${\mathrm{A}}$ to its associated graded object $({\mathrm{A}}_{{\mathrm{i}}}/{{\mathrm{A}}_{{\mathrm{i}}-1}})_{{\mathrm{i}}\in{\mathbb{Z}}}$. The functor $\gamma$ restricts to a functor
$\gamma^{\prime}:\mathrm{Fun}({\mathbb{N}},{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\to\prod_{{\mathbb{N}}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}.$
Let $\theta$ be the functor
$\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to\mathrm{Fun}({\mathbb{N}},\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})),\ {\mathrm{A}}\mapsto(...\to{\mathcal{O}}\circ_{\mathsf{f}_{{\mathrm{n}}}({\mathcal{O}})}{\mathrm{A}}\to{\mathcal{O}}\circ_{\mathsf{f}_{{\mathrm{n}}+1}({\mathcal{O}})}{\mathrm{A}}\to...).$
@@
Moreover by [21, Remark 4.23.] the composition
[…]
factors as the forgetful functor $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathcal{C}}$ followed by the trivial Lie algebra functor ${\mathcal{C}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}).$ Thus the functor ${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ factors as
[…]
$\xrightarrow{\mathrm{colim}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
[…]
$\xrightarrow{\gamma^{\prime}}\prod_{{\mathbb{N}}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\xrightarrow{\oplus}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
$\mathrm{Sym}:{\mathcal{C}}\xrightarrow{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}\circ(-)}{\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$

### v2

@@ 3. Koszul duality
$\tau_{n}({\mathcal{O}})*_{{\mathcal{O}}}{\mathrm{triv}}\simeq\tau_{n}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}}.$
Construction 3.11.
@@
[…]
Recall that by [39, Example 6.1.6.22.] for every object $\mathrm{X}$} of $\mathcal{C}$} equipped with an action of a finite group $\mathrm{G}$} there is a norm map ${\mathrm{X}}_{\mathrm{G}}\to{\mathrm{X}}^{\mathrm{G}}$ from the coinvariants to the invariants, which we will use in the following.
@@
Let $\mathcal{C}$} be a stable symmetric monoidal $\infty$-category compatible with small colimits that admits small limits, $\mathcal{O}$} a non-unital $\infty$-operad in $\mathcal{C}$} such that ${\mathcal{O}}_{1}={\mathbb{1}}$ and $\mathrm{X}$} an $\mathcal{O}$}-algebra. If all norm maps associated to symmetric groups are equivalences, for every ${n}\in{\mathbb{N}}$ the following morphism is an equivalence:
$\rho:\tau_{n}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{X}}\to(\tau_{n}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}({\mathrm{triv}}\circ_{{\mathcal{O}}}{\mathrm{X}}).$
@@
We proceed by induction on ${n}\geq 1.$ For n=1{n}=1 the map $\rho$ identifies with the canonical equivalence
@@
The fiber $\mathrm{T}$} of the map $\tau_{{n}+1}({\mathcal{O}})\to\tau_{n}({\mathcal{O}})$ of right $\mathcal{O}$}-modules is the symmetric sequence concentrated in degree n+1{n}+1 with value ${\mathcal{O}}_{{n}+1}$ equipped with the trivial right $\mathcal{O}$}-module structure according to Lemma 2.46. Applying the exact functor $(-)\circ_{\mathcal{O}}{\mathrm{X}}:{\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}}))\to{\mathrm{sSeq}}({\mathcal{C}})$ we obtain the fiber sequence
${\mathrm{T}}\circ{\mathrm{Y}}={\mathrm{T}}\circ({\mathrm{triv}}\circ_{\mathcal{O}}{\mathrm{X}})\to\tau_{{n}+1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}\to\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{X}}.$
@@
$(\tau_{{n}+1}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}){*}^{{\mathcal{O}}^{\vee}}{\mathrm{Y}}\to(\tau_{n}({\mathcal{O}})\circ_{\mathcal{O}}{\mathrm{triv}}){*}^{{\mathcal{O}}^{\vee}}{\mathrm{Y}}$
@@
[…]
where both horizontal sequences are fiber sequences. The left vertical map ${\mathrm{T}}\circ{\mathrm{Y}}\to{\mathrm{T}}{\ast}{\mathrm{Y}}$ identifies with the norm map $({\mathcal{O}}_{{n}+1}\otimes{\mathrm{Y}}^{\otimes{n}+1})_{\Sigma_{{n}+1}}\to({\mathcal{O}}_{{n}+1}\otimes{\mathrm{Y}}^{\otimes{n}+1})^{\Sigma_{{n}+1}}.$
@@
[…]
@@
[…]
@@
Recently, Heuts [31] generalizes Theorem 3.13 to the situation that norm maps are not necessarily invertible.
@@ 4. A derived Milnor-Moore theorem
Computing the Koszul dual one finds that for every ${n}\in\Sigma$ there is a canonical equivalence of spectra
${\mathrm{Lie}}_{n}\simeq\bigoplus_{{n}-1!}{\mathbb{1}}_{\mathrm{Sp}},$
while the $\Sigma_{n}$-action on ${\mathrm{Lie}}_{n}$ is highly non-trivial. In particular, $\mathrm{Lie}$} is an $\infty$-operad in connective spectra ${\mathrm{Sp}}_{\geq 0}.$
@@
${\mathrm{Lie}}_{\mathcal{C}}=\alpha^{\mathcal{C}}_{!}({\mathrm{Lie}})\simeq({\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}})^{\vee}(-1).$
@@
    $\theta:\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
    ${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})$
@@
1. We like to see that for any ${\mathrm{X}},{\mathrm{Y}}\in\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})$ the canonical map $\theta({\mathrm{X}}\times{\mathrm{Y}})\to\theta({\mathrm{X}})\times\theta({\mathrm{Y}})$ is an equivalence. By Lemma 2.30 the forgetful functors $\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\to{\mathcal{C}}$ and $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\to{\mathcal{C}}$ are comonadic. Let ${\mathrm{Sym}}_{>0}$ be the right adjoint of the forgetful functor $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\to{\mathcal{C}}$. By comonadicity any ${\mathrm{X}}\in\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})$ is the totalization of a cosimplicial object $\mathrm{X}$$^{\prime}$ in $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})$ taking values in the essential image of ${\mathrm{Sym}}_{>0}$, whose image in $\mathcal{C}$} splits. Hence $\theta$ preserves the limit of $\mathrm{X}$$^{\prime}$ and the canonical map $\theta({\mathrm{X}}\times{\mathrm{Y}})\to\theta({\mathrm{X}})\times\theta({\mathrm{Y}})$ identifies with the map
@@
${\mathrm{Sym}}:{\mathcal{C}}\xrightarrow{{\mathrm{Sym}}_{>0}}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\xrightarrow{\theta}\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
preserves finite products. By [39, Proposition 3.2.4.7.] the object-wise symmetric monoidal structure on $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ is cartesian. For any ${\mathrm{X}},{\mathrm{Y}}\in{\mathcal{C}}$ the canonical map ${\mathrm{Sym}}({\mathrm{X}}\oplus{\mathrm{Y}})\to{\mathrm{Sym}}({\mathrm{X}})\otimes{\mathrm{Sym}}({\mathrm{Y}})$ identifies with the canonical equivalence in $\mathcal{C}$}.
2.: In view of (1) it is enough to see that the composition
$\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-)}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
preserves finite products.
In the following we consider $\mathbb{N}$$\mathbb{Z}$} as categories by viewing them as posets. Let
$\gamma:\mathrm{Fun}({\mathbb{Z}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\to\prod_{{\mathbb{Z}}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
be the functor that sends a filtered object $\mathrm{A}$} to its associated graded object $({\mathrm{A}}_{{i}}/{{\mathrm{A}}_{{i}-1}})_{{i}\in{\mathbb{Z}}}$. The functor $\gamma$ restricts to a functor
$\gamma^{\prime}:\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\to\prod_{{\mathbb{N}}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}.$
Let $\theta$ be the functor
@@
Moreover by [30, Remark 4.23.] the composition
$\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{\theta}\mathrm{Fun}({\mathbb{N}},\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{\gamma^{\prime}}\prod_{{\mathbb{N}}}\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{\oplus}\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})$
factors as the forgetful functor $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathcal{C}}$ followed by the trivial Lie algebra functor ${\mathcal{C}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}).$ Thus the functor ${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ factors as
$\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{\theta}\mathrm{Fun}({\mathbb{N}},\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))_{*}}\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})$
$\xrightarrow{\mathrm{colim}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
$\rho:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{\theta}\mathrm{Fun}({\mathbb{N}},\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))_{*}}\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})$
$\xrightarrow{\gamma^{\prime}}\prod_{{\mathbb{N}}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\xrightarrow{\oplus}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
${\mathrm{Sym}}:{\mathcal{C}}\xrightarrow{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}\circ(-)}\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$